OpenStudy (anonymous):

The ordered pairs (1, 25), (2, 36), (3, 49), (4, 64), and (5, 81) represent a function. What is a rule that represents this function? (1 point) y = (5x)^2 y = x + 11 y = x^5 y = (x + 4)2

OpenStudy (amistre64):

input the x values to see what produces the y values for the given points

OpenStudy (anonymous):

This is similar to the last one you just did... you just try the x values to see. But if you have some idea of which one looks promising, try that one first, or if you can tell immediately that one is wrong, don't bother trying all the points.

OpenStudy (anonymous):

(x+4)^2?

OpenStudy (amistre64):

what do we get for y values when we use that one?

OpenStudy (anonymous):

I just. I don't even know I am so confused.

OpenStudy (anonymous):

you can make a table to try the values... Here are the points (1, 25), (2, 36), (3, 49), (4, 64), and (5, 81). Put the x value and y value for each point in the first two columns, then in the third column, calculate what the rule says should be the y value. I will do the first one... x y rule using "x", y = (x+4)^2 --------------------------------- 1 25 (1+4)^2 = 25 <<--- rule is correct for this point 2 36 (2+4)^2 =

OpenStudy (anonymous):

I think you probably weren't too confused, but something in this discussion made you doubt yourself.

OpenStudy (campbell_st):

the input into the is 1 2, 3, 4, 5.... the output from the function is 25, 36, 49, 64, 81 which are all square numbers... so thats the clue... your choice was correct... just check by substituting x = 1 to make sure you get f(1) = 25, and x =3 to fet f(3) = 49

OpenStudy (anonymous):

Great point :) So @lexxiepoopoo , now you have two "clues" to consider... when the y values get huge fast, then you probably have an "x" in your exponent, like the last problem. And if your y values are square numbers, then you have a squared function of some sort... Using clues like that helps you not waste time trying out every single rule...